40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 14 16:20:11 2011, Zach, Update, SUS, C1:SUS-ETMX_SPDMon fixed board_works.pngrut_roh_raggie.pngScreen_shot_2011-10-14_at_4.09.45_PM.png
    Reply  Fri Oct 14 18:38:41 2011, kiwamu, Update, SUS, C1:SUS-ETMX_SPDMon fixed 
Message ID: 5666     Entry time: Fri Oct 14 16:20:11 2011     Reply to this: 5667
Author: Zach 
Type: Update 
Category: SUS 
Subject: C1:SUS-ETMX_SPDMon fixed 

I offered to help Kiwamu with some of the 40m work. The first task was to figure out why the ETMX side OSEM monitor was so low, since we know that the depth is about right. It was showing ~0.13 V to the others' ~0.7 V.

TL,WR: There was a wire disconnected from the breakout panel on the side of the rack

I started by pulling the board out and checking to make sure that it was working properly. I injected a sine wave to the SIDE IN and found that it showed up in the signal coming out of the back (into the crate) just fine (see below). One strange thing I noticed while testing the board is that both inputs for each used channel of the MAX333 switches on the board are shorted to their respective outputs. That is, the switches seem to be open to BOTH 0 and 1 logic states. This seems counterintuitive, but perhaps there's something about how these work that I don't know.



Then I went about tracing the signal from the back of the crate to the breakout panel on the side of the rack. I opened it up, verified that the ribbon cable was transmitting correctly, and as I went to plug it back in I noticed that one of the wires---the correct one---had come completely undone.


The screw clamp appeared to be a bit finicky, as I had to loosen and tighten it a few times before it finally seemed to grab hold of the wire. It probably just wasn't tight in the first place and the wire was pulled out. Anyway, things are working now:


ELOG V3.1.3-